# On the Learnability of Watermarks for Language Models

### v2

@@ 3.3 Sampling-based watermark distillation
Sampling-based watermark distillation (WD-S) consists of two stages. First, we generate watermarked text from teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ applied using key $\xi$. Then, we fine-tune the student model $p_{\theta}$ on this watermarked text using the standard language modeling cross-entropy loss.
Formally, given a prompt $x$, we generate a watermarked completion sequence $y=y_{n}y_{n-1}\cdots y_{1}$ where $y_{t}\sim f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid y_{<t}x\right),y_{<t}x,\xi\right)$. Given a set of prompts $\mathcal{P}$, we populate a finetuning dataset $\mathcal{D}$ with watermarked sequences $y$ from prompts $x\in\mathcal{P}$. Then, we train $p_{\theta}$ to minimize the cross-entropy loss
$\mathcal{L}_{\text{WD-S}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{y\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(y)}-\log p_{\theta}\left(y_{t}\mid y_{<t}\right).$ (9)
Here, $p_{\mathrm{LM}}$ and $p_{\theta}$ do not need to share the same tokenizer or vocabulary. However, sampling-based watermark distillation is less efficient than logits-based watermark distillation due to autoregressively generating watermarked text.
@@ 4 Experimental setup
4.1 Watermarking strategies
We experiment with all three decoding-based watermarking strategies discussed in §2. We use various hyperparameter settings to vary the level of distortion induced by the watermarks. Specifically, we test KGW with $\gamma=0.25$ and $\delta=\{1,2\}$, [^5] Because we always use $\gamma=0.25$, we sometimes omit explicitly stating the value of $\gamma$ to reduce clutter. Aar with $k=\{2,3,4\}$, and KTH with key length 256 and number of shifts $s=\{1,2,4,256\}$.
@@ 4.2 Training
For each watermarking strategy, we test logits-based and sampling-based watermark distillation.
For logits-based watermark distillation, we use Llama 2 7B (Touvron et al., 2023) as both the teacher and student models (the student model is initialized with the teacher model weights). We distill using a subset of OpenWebText (Gokaslan et al., 2019) for 5,000 steps with a batch size of 64, sequence length of 512 tokens, [^6] For KTH we use a batch size of 128 and sequence length of 256 tokens because we use key length 256. maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix A.1.
For sampling-based watermark distillation, we use Llama 2 7B as the teacher model and Pythia 1.4B (Biderman et al., 2023) as the student model. [^7] We include additional experiments using Llama 2 7B as both the teacher and student model for sampling-based watermark distillation in Appendix C. Note that Llama 2 and Pythia have different tokenizers, so sampling-based distillation is necessary for this setting. First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640k watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. [^8] We include additional experiments using fewer training samples to investigate sample efficiency in Appendix D. Then, we fine-tune Pythia 1.4B on the watermarked samples for 1 epoch, roughly 8,000 steps, with a batch size of 64, sequence length of 256 tokens, [^9] Note that Llama 2 and Pythia tokenize sequences differently. Pythia has a larger vocabulary size and tends to tokenize sequences into fewer tokens compared to Llama 2. maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix A.2.
@@ 4.3 Evaluation and metrics
Evaluation procedure. As in Kirchenbauer et al. (2023a) and Kuditipudi et al. (2023), we evaluate on generations prompted by prefixes from the RealNewsLike subset of the C4 dataset (Raffel et al., 2020). For each decoding-based watermarking strategy and distilled model, we generate 5k 200-token completions from 50-token prompts from the validation split. We use standard sampling with temperature 1 for the main results, and investigate the model’s robustness to different decoding parameters in §5.3. We include evaluations on additional datasets in Appendix F.
@@
Watermark detectability. We compute the median watermark detection p-value across generations. [^10] For watermark detection, we truncate to the first 200 tokens using the detection tokenizer (Llama 2). Note that the p-values for the KTH watermark are lower bounded by how many samples $T$ we compute in the reference distribution. Similar to Kuditipudi et al. (2023), we use $T=10,000$, [^11] As in Kuditipudi et al. (2023), we do not use higher $T$ due to the slowness of computing the test statistic. so the p-values are lower bounded by 1e-04. So, to make fine-grained distinctions in watermark strength below this lower bound, we also compute the median test statistic (defined in §2.3) to evaluate KTH watermark strength. A larger test statistic indicates higher watermark detectability.
@@
Generation quality. We compute the mean LM-score across generations, which we define to be perplexity evaluated by Llama 2 13B. [^12] Note that LM-score does not perfectly capture quality or fluency, as repetitive text has a low LM-score, so we also use separate metrics measuring repetitiveness. We also compute MAUVE to measure distributional similarity between model generations and human text (Pillutla et al., 2021). To evaluate repetitiveness, we compute mean seq-rep-3 across generations, which is the proportion of duplicate 3-grams in a sequence (Welleck et al., 2020).
Comparisons and notation. For both watermark distillation experiments, for each decoding-based watermarking strategy $f_{\text{w}}$, we compare the teacher model with $f_{\text{w}}$ (denoted by DecT) against the distilled student model (denoted by WD-L and WD-S for logits-based and sampling-based watermark distillation, respectively). As a baseline for generation quality, we use the base student model with no watermarking or distillation (denoted by “Base student”).
For the sampling-based watermark distillation experiment, since the teacher Llama 2 7B model is larger and more powerful than the student Pythia 1.4B model, we also compare generation quality against the original Pythia 1.4B model with $f_{\text{w}}$ (denoted by DecS) to control for model size. This allows for a fair comparison of generation quality between decoding-based watermarking and sampling-based watermark distillation.
@@ 5 Results
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 4.4e-18 6.9e-09 1.000 0.997 16.4 17.5 0.04 0.03 0.89 0.90
$\delta=1$ 1.0e-05 6.0e-03 0.992 0.913 12.6 13.1 0.03 0.03 0.91 0.90
Aar $k=2$ 1.3e-75 2.2e-12 0.999 0.996 6.5 10.8 0.34 0.12 0.28 0.83
$k=3$ 4.8e-73 1.1e-01 1.000 0.773 9.5 11.5 0.14 0.04 0.74 0.93
$k=4$ 4.0e-72 3.7e-01 0.999 0.580 10.7 11.8 0.09 0.03 0.86 0.92
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (569) 0.994 0.998 10.5 16.6 0.03 0.04 0.88 0.83
$s=2$ 1.0e-04 (600) 1.0e-04 (478) 0.998 0.994 10.7 16.5 0.03 0.04 0.91 0.84
$s=4$ 1.0e-04 (593) 1.0e-03 (439) 0.995 0.969 10.5 14.0 0.03 0.04 0.91 0.90
$s=256$ 1.0e-04 (600) 4.6e-02 (426) 0.996 0.893 10.6 11.4 0.03 0.04 0.92 0.92
Base student 5.0e-01 0.500 11.7 0.03 0.92
@@
Table 1: Results for the logits-based watermark distillation experiments. Within each watermark type, the hyperparameter rows go from higher-distortion to lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently.
@@
Figure 2: Empirical cumulative distribution functions (eCDFs) of the distributions of watermark detection p-values on generations from logits-based (left) and sampling-based (right) watermark distilled models on various watermark types and hyperparameter settings. In higher-distortion watermarks, the majority of generations have small p-values. In lower-distortion watermarks, the p-values are larger, but still consistently smaller than a non-watermarked baseline.
5.1 Logits-based watermark distillation
Table 1 contains results for logits-based watermark distillation. [^13] As discussed in §4.3, we report the median KTH test statistic in parentheses to allow for finer-grained distinctions in watermark strength below the p-value lower bound of 1.0e-04. We find that logits-based watermark distillation (WD-L) successfully learns higher-distortion [^14] Here, we are using “distortion” somewhat informally, roughly meaning how much of a difference watermarking causes in terms of generation quality, behavior, etc. watermarks, achieving small p-values and high detectability. Lower-distortion watermarks are harder to learn, but they are still learnable, just less efficiently. Table 1 shows that WD-L does not achieve as small watermark detection p-values as decoding-based watermarking, but for higher-distortion watermark hyperparameter settings (larger $\delta$ for KGW, smaller $k$ for Aar, and smaller $s$ for KTH), the p-values are still sufficiently small to enable high detectability, as shown by the high AUROC scores. Figure 2 (left) contains empirical CDFs of the distributions of p-values across generations, showing that for higher-distortion watermarks, the majority of generations from WD-L have small p-values.
Within each watermark type, WD-L p-values are larger for lower-distortion hyperparameter settings, indicating that lower-distortion watermarks are harder to learn. These watermarks are still learned to some degree, however, as the p-values are noticeably smaller than the random chance baseline of 0.5, and the AUROC scores are noticeably higher than the baseline of 0.5. We also find that if we run logits-based watermark distillation on Aar $k=3$ for five times longer (25,000 steps) on more data, the median p-value decreases from 0.11 to 0.00014. This suggests that lower-distortion watermarks are less sample efficient to learn, but still learnable, given sufficient training data and steps.
In addition, we find that for each watermark type, WD-L weights-based and decoding-based watermarking distort generation quality along the same axes, although the degree of distortion varies. For example, for Aar, both WD-L and decoding-based watermarking are more repetitive than base Llama 2 7B, especially at lower $k$. However, WD-L induces less distortion than decoding-based Aar, as shown by the lower repetition and higher MAUVE scores of WD-L. On the other hand, WD-L induces more distortion than decoding-based watermarking on KGW and KTH, as shown by the higher LM-scores of WD-L.
5.2 Sampling-based watermark distillation
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 4.4e-18 3.6e-06 1.000 0.995 16.4 32.8 51.9 0.89 0.81 0.65
$\delta=1$ 1.0e-04 2.0e-02 0.992 0.874 12.6 26.2 41.3 0.91 0.84 0.72
Aar $k=2$ 1.3e-75 2.9e-14 0.999 0.993 6.5 7.0 19.8 0.28 0.10 0.52
$k=3$ 4.8e-73 1.6e-02 1.000 0.866 9.5 15.1 29.0 0.74 0.46 0.68
$k=4$ 4.0e-72 3.0e-01 0.999 0.631 10.7 20.5 33.8 0.86 0.76 0.75
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (456) 0.994 0.989 10.5 22.2 34.0 0.88 0.72 0.81
$s=2$ 1.0e-04 (600) 4.0e-04 (445) 0.998 0.972 10.7 21.9 26.3 0.91 0.79 0.84
$s=4$ 1.0e-04 (593) 2.2e-03 (436) 0.995 0.955 10.5 22.2 24.7 0.91 0.84 0.85
$s=256$ 1.0e-04 (600) 2.1e-03 (437) 0.996 0.955 10.6 22.2 25.7 0.92 0.85 0.88
Base student 5.0e-01 0.500 21.7 0.88
Table 2: Results for the sampling-based watermark distillation experiments. Within each watermark type, the hyperparameters become lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently.
Table 2 contains results for the sampling-based watermark distillation experiments. [^15] As discussed in §4.3, we report the median KTH test statistic in parentheses to allow for finer-grained distinctions in watermark strength below the p-value lower bound of 1.0e-04.161616The seq-rep-3 metric is deferred to Appendix B due to space constraints. The trend is the same as in logits-based watermark distillation, with Aar increasing repetitiveness across all methods, especially at low $k$. We find the same trends and conclusions as in the logits-based watermark distillation experiments. Table 2 and Figure 2 (right) show that sampling-based watermark distillation (WD-S) successfully learns higher-distortion watermarks, achieving small p-values and higher detectability. The p-values from WD-S are not as small as those from decoding-based watermarking, but still small enough to enable high detectability, as shown by the high AUROC scores. Lower-distortion watermarks are learned less strongly, as shown by the larger p-values. However, in sample complexity experiments in Appendix D, we find that more training samples and steps lead to smaller p-values, with no sign of convergence at 640k samples, which is how many we use here. This suggests that lower-distortion watermarks are less sample efficient to learn, but still learnable, given enough training samples.
5.3 Robustness to changes in decoding algorithm

### v3

@@ 3.2 Sampling-based watermark distillation
Sampling-based watermark distillation has two stages. First, we generate watermarked text from teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ applied using key $\xi$. Then, we fine-tune the student model $p_{\theta}$ on this watermarked text using the standard language modeling cross-entropy loss.
Formally, given a set of prompts $\mathcal{P}$, for each prompt $z\in\mathcal{P}$, we generate a watermarked completion sequence $x=x_{1}x_{2}\cdots x_{n}$, where each sampled token $x_{t}\sim f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid zx_{<t}\right),zx_{<t},\xi\right)$. Let the fine-tuning dataset $\mathcal{D}$ consist of these watermarked generations $x$ [^5] Ideally, the intended use case and domain of the student model $p_{\theta}$ should inform the choices of the set of prompts $\mathcal{P}$ and teacher model $p_{\mathrm{LM}}$. However, empirically, we find that sampling-based watermark distillation is fairly robust to domain shifts (see §4.3 and Appendix F). Then, we train $p_{\theta}$ to minimize the cross-entropy loss on $\mathcal{D}$, given by
$\mathcal{L}_{\mathrm{sampling}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(x)}-\log p_{\theta}\left(x_{t}\mid x_{<t}\right).$ (4)
Here, $p_{\mathrm{LM}}$ and $p_{\theta}$ do not need to share the same tokenizer or vocabulary. However, sampling-based watermark distillation is less efficient than logit-based watermark distillation due to autoregressively generating watermarked text in the first stage.
@@ 4 Experimental setup
We run experiments to evaluate how well logit-based and sampling-based watermark distillation can learn weights-based watermarking from the decoding-based watermarking strategies seen in §2. Ideally, we want weights-based watermarking to match decoding-based watermarking in terms of watermark detectability and generation quality.
4.1 Watermarking strategies and hyperparameters
We experiment with the three decoding-based watermarking strategies discussed in §2. We use various hyperparameter settings to vary the level of distortion induced by the watermarks. Specifically, we test KGW with $k=\{0,1,2\}$, $\gamma=0.25$ and $\delta=\{1,2\}$, [^6] Because we always use $\gamma=0.25$, we sometimes omit explicitly stating the value of $\gamma$ to simplify notation. [^7] We exclude $k=2,\delta=1$ since we find that $k=2,\delta=2$ already exhibits lower learnability. Aar with $k=\{2,3,4\}$, and KTH with key sequence length $m=256$ and number of shifts $s=\{1,2,4,256\}$.
@@ 4.2 Training
For each decoding-based watermarking strategy, we test logit-based and sampling-based watermark distillation for learning weights-based watermarking.
For logit-based watermark distillation, we use Llama 2 7B (Touvron et al., 2023) as both the teacher and student models (the student model is initialized with the teacher model weights). We distill using a subset of OpenWebText (Gokaslan et al., 2019) for 5,000 steps with a batch size of 64 sequences, sequence length of 512 tokens, [^8] For KTH we use a batch size of 128 and sequence length of 256 tokens since we use key length $m=256$. maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix B.1.
For sampling-based watermark distillation, we also use Llama 2 7B as both the teacher and student models. First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640,000 watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. Then, we fine-tune Llama 2 7B on the watermarked samples for 1 epoch of 5,000 steps, with a batch size of 128 sequences, sequence length of 256 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix B.2.
In Appendix C, we perform sampling-based watermark distillation experiments where the teacher and student models have different tokenizers and sizes, using Llama 2 7B as the teacher model and Pythia 1.4B as the student model (Biderman et al., 2023).
@@ 4.3 Evaluation and metrics
Evaluation procedure. As in Kirchenbauer et al. (2023a) and Kuditipudi et al. (2023), we evaluate on generations prompted by prefixes from the RealNewsLike subset of the C4 dataset (Raffel et al., 2020). For each decoding-based watermarking strategy and distilled model, we generate 5,000 200-token completions from 50-token prompts from the validation split. We use standard sampling with temperature 1 for the main results, and investigate the model’s robustness to different decoding parameters in §5.1. We include evaluations on additional datasets in Appendix F.
@@
Watermark detectability. We compute the median watermark detection p-value across generations. Note that the p-values for the KTH watermark are lower bounded by how many samples $T$ we compute in the reference distribution. Similar to Kuditipudi et al. (2023), we use $T=$10,000$$, so the p-values are lower bounded by 1e-4. To make finer-grained distinctions in watermark strength below this lower bound, we also compute the median test statistic (discussed in §2.3) to evaluate KTH watermark strength. A lower (more negative) test statistic indicates higher watermark detectability.
@@
Generation quality. We use Llama 2 13B to compute the mean perplexity of generations. Lower perplexity tends to indicate higher quality and fluency, but repetitive text also achieves low perplexity. So, to evaluate repetitiveness, we compute the mean seq-rep-3 of generations, which is the proportion of duplicate 3-grams in a sequence, given by 1# of unique 3-grams# of $1-\frac{\text{\# of unique 3-grams}}{\text{\# of 3-grams}}$ (Welleck et al., 2020).
Comparisons. For both watermark distillation methods, for each decoding-based watermarking strategy $f_{\text{w}}$, we compare the teacher model with $f_{\text{w}}$ applied (denoted by “Decoding”) against the distilled student model (denoted by “Logit” and “Sampling” for logit-based and sampling-based watermark distillation, respectively). As a baseline for generation quality, we use the base student model with no watermarking or distillation (denoted by “Base student”).
@@ 5 Results
Table 1 contains results for the logit-based and sampling-based watermark distillation experiments. The two watermark distillation methods exhibit similar trends. Both forms of watermark distillation successfully learn higher-distortion watermarks, [^9] Here, we are using “distortion” somewhat informally, roughly meaning how much of a difference watermarking causes in terms of generation quality, behavior, etc. achieving small p-values and high detectability. In some watermarks, e.g., KGW $k=0$, watermark distillation matches the p-values achieved by decoding-based watermarking. In other watermarks, watermark distillation does not achieve as small watermark detection p-values as decoding-based watermarking, but for higher-distortion watermark hyperparameter settings (smaller $k$ and larger $\delta$ for KGW, smaller $k$ for Aar, and smaller $s$ for KTH), the p-values are still sufficiently small to enable high detectability, as shown by the high AUROC values. Figure 2 contains empirical CDFs of the distributions of p-values across generations, showing that for higher-distortion watermarks, the majority of generations from the watermark distilled models have small p-values.
Within each watermark type, p-values from logit-based and sampling-based distillation are larger for lower-distortion hyperparameter settings, indicating that lower-distortion watermarks are harder to learn. However, these watermarks are still learned to some degree, as the p-values are noticeably smaller than the non-watermarked baseline of 0.5, and the AUROC values are noticeably higher than the non-watermarked baseline of 0.5. In Appendix D, sample complexity experiments show that more training samples and steps lead to smaller p-values for both logit-based and sampling-based distillation, with no sign of convergence. In addition, we find that when we train logit-based watermark distillation on KGW $k=2,\delta=2$ for five times longer (25,000 steps) on more data, the median p-value decreases from 0.1 to 0.012. This suggests that lower-distortion watermarks are less sample efficient to learn, but still learnable, given sufficient training data and steps.
Compared to decoding-based watermarking, watermark distillation does not achieve as optimal a tradeoff between generation quality and detectability. For KGW and KTH, both watermark distillation methods achieve slightly to moderately higher perplexity and similar or larger p-values compared to decoding-based watermarking. For Aar, watermark distillation achieves similar or lower seq-rep-3 as decoding-based watermarking, but larger p-values. This suggests that to learn weights-based watermarking, logit-based and sampling-based watermark distillation incur some cost to the tradeoff between generation quality and detectability.
While logit-based and sampling-based watermark distillation show similar trends, there are some interesting differences. On the Aar watermark, sampling-based distillation and decoding-based watermarking have similarly high repetitiveness, whereas logit-based distillation achieves significantly less repetitiveness. We hypothesize that this is because sampling-based distillation trains on entire repetitive sequences generated by decoding-based watermarking, whereas logit-based distillation trains only on next-token predictions on non-repetitive human-generated prefixes. Also, on the KTH watermark, sampling-based distillation achieves higher detectability and lower perplexity than logit-based distillation, particularly at larger numbers of shifts $s$. We speculate that this may be because sampling-based distillation trains on complete sequences that are each watermarked with a consistent shift $\tau$, whereas logit-based distillation trains only on next-token predictions on non-watermarked prefixes that contain no shift information. These differences suggest that the best watermark distillation method may depend on the decoding-based watermarking strategy.
However, recall that logit-based and sampling-based distillation have different requirements (e.g., access to logits and shared tokenizer vs. access to samples and autoregressive generation, see §3.1 and §3.2), so they should not be compared solely on performance. So, logit-based and sampling-based distillation are each suitable and applicable for different settings, so neither is strictly better than the other in all scenarios.
@@
Watermark
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
Decoding
Logit
Sampling
KGW $k=0,\delta=2$ 6e-16 2e-17 2e-15 1.00 1.00 1.00 17.5 17.3 20.3 0.05 0.05 0.05
$k=1,\delta=2$ 4e-18 7e-09 8e-07 1.00 1.00 1.00 16.5 17.6 19.2 0.04 0.03 0.04
$k=2,\delta=2$ 9e-18 1e-01 1e-01 1.00 0.80 0.74 16.8 17.7 19.8 0.03 0.02 0.03
$k=0,\delta=1$ 5e-04 3e-05 1e-03 0.98 0.99 0.98 13.0 12.9 15.7 0.03 0.03 0.03
$k=1,\delta=1$ 1e-05 7e-03 2e-02 0.99 0.91 0.87 12.7 13.1 14.9 0.03 0.03 0.03
Aar $k=2$ 1e-75 2e-12 3e-17 1.00 1.00 0.98 6.5 10.8 7.7 0.34 0.11 0.34
$k=3$ 5e-73 1e-01 6e-03 1.00 0.78 0.88 9.5 11.6 10.5 0.14 0.04 0.17
$k=4$ 4e-72 4e-01 3e-01 1.00 0.58 0.65 10.7 11.8 11.9 0.09 0.03 0.11
KTH $s=1$ 1e-04 (-593) 1e-04 (-565) 1e-04 (-561) 1.00 1.00 1.00 10.5 16.5 15.1 0.03 0.04 0.03
$s=2$ 1e-04 (-596) 1e-04 (-476) 1e-04 (-525) 1.00 0.99 0.99 10.7 16.3 13.4 0.03 0.04 0.03
$s=4$ 1e-04 (-594) 1e-03 (-438) 1e-04 (-487) 1.00 0.96 0.99 10.6 14.2 12.5 0.03 0.04 0.04
$s=256$ 1e-04 (-594) 8e-02 (-423) 1e-04 (-453) 1.00 0.85 0.97 10.8 11.3 11.3 0.03 0.04 0.04
Base student 5e-01 0.50 11.8 0.03
@@
Table 1: Results for logit-based and sampling-based watermark distillation experiments. Within each watermark type (KGW, Aar, and KTH), the hyperparameter rows go from higher-distortion to lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently and strongly.
(a) Logit-based distillation eCDFs
(b) Sampling-based distillation eCDFs
@@
Figure 2: Empirical cumulative distribution functions (eCDFs) of watermark detection p-values of generations from logit-based (a) and sampling-based (b) watermark distillation. In higher-distortion watermarks, the majority of generations have small p-values. In lower-distortion watermarks, the p-values are larger, but still consistently smaller than a non-watermarked uniform baseline.
5.1 Robustness to changes in decoding parameters
Whereas decoding-based watermarking relies on specialized decoding algorithms, weights-based watermarking generates watermarked text naturally under standard decoding algorithms. Table 2 shows watermark detection p-values from weights-based watermarking learned by logit-based and sampling-based watermark distillation under a variety of decoding algorithms and parameters: nucleus sampling (Holtzman et al., 2020) with different thresholds $p$, temperature-based sampling with different temperatures $t$, and greedy decoding ($t=0$). All settings achieve consistently small p-values, showing that weights-based watermarking is robust to changes in decoding parameters. The p-values decrease as $t$ or $p$ decreases, with greedy decoding achieving the smallest p-values.